OpenStudy (jazzyluvsu123):

lance rode a total of 46&3/8 on Saturday. He rode 19&3/4 miles in the morning. How far did he ride in after noon? Giving medal for however I think explains the best @magix430 @Directrix

OpenStudy (anonymous):

What they are asking you to do is subtract the second number from the first. What I would do is make them both into improper fractions, find a common denominator, do the subtraction, and then simplify.

OpenStudy (jazzyluvsu123):

ok let me try it on paper then I will draw it out and see you can see if it is correct!

OpenStudy (anonymous):

\[46\frac{ 3 }{ 8 } - 19\frac{ 3 }{ 4 } = x\] Turn them all into improper fractions: \[\frac{ 371 }{ 8 } - \frac{ 79 }{ 4 }\] Make the bottoms of each into 8 (or the common denominator). To do so, multiply the fraction with denominator 4 by 2. \[\frac{ 371 }{ 8 } - \frac{ 158 }{ 8 }\] Subtract as usual: \[\frac{ 213 }{ 8 }\] Simplify: - It cannot be simplified.

OpenStudy (anonymous):

Revert back into proper fraction, that is: \[26\frac{ 5 }{ 8 }\]

OpenStudy (jazzyluvsu123):

ok thank u